OpenStudy (anonymous):

I have no idea how to do this could someone explain this to me If f(x)=x^2+3x+5 What is f(a+h)

OpenStudy (anonymous):

http://imgur.com/ErkWWdf

OpenStudy (anonymous):

those are the choices

OpenStudy (bradely):

f(x)=x^2+3x+5 x replace with a+h f(a+h)=(a+h)^2+3(a+h)+5 =a^2+2ah+h^2+3a+3h+5 so option B correct Further any help post on http://www.mathskey.com/question2answer/ask

OpenStudy (anonymous):

@bradely thanks a lot i kind of understand it now